# Quantize What Counts: Bit Allocation Insights Informed by Spectral Gaps in Keys and Values

### v2

@@ Results.
Qwen3-0.6B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+3\%}}{{<}}\mathbf{0.03}$ $0.00\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{<}}\mathbf{0.16}$ $0.04\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{<}}\mathbf{0.17}$ $\mathbf{0.34}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.18$ $\mathbf{0.34}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.18$
Llama-3.2-1B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+5\%}}{{<}}\mathbf{0.05}$ $0.01\stackrel{{\scriptstyle\scriptscriptstyle+18\%}}{{<}}\mathbf{0.19}$ $\mathbf{0.27}\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{>}}0.20$ $\mathbf{0.28}\stackrel{{\scriptstyle\scriptscriptstyle+8\%}}{{>}}0.20$ $\mathbf{0.28}\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{>}}0.21$
@@
Qwen3-4B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+41\%}}{{<}}\mathbf{0.41}$ $0.01\stackrel{{\scriptstyle\scriptscriptstyle+49\%}}{{<}}\mathbf{0.50}$ $\mathbf{0.80}\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{>}}0.51$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.51$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.51$
@@
Qwen3-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+48\%}}{{<}}\mathbf{0.48}$ $0.08\stackrel{{\scriptstyle\scriptscriptstyle+46\%}}{{<}}\mathbf{0.54}$ $\mathbf{0.86}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.55$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+32\%}}{{>}}0.55$ $\mathbf{0.86}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.55$
Qwen3-14B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+50\%}}{{<}}\mathbf{0.50}$ $0.26\stackrel{{\scriptstyle\scriptscriptstyle+32\%}}{{<}}\mathbf{0.58}$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}0.60$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}0.60$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}0.60$
@@
Qwen3-0.6B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+10\%}}{{<}}\mathbf{10.06}$ $0.00\stackrel{{\scriptstyle\scriptscriptstyle+53\%}}{{<}}\mathbf{52.51}$ $51.23\stackrel{{\scriptstyle\scriptscriptstyle+2\%}}{{<}}\mathbf{52.87}$ $\mathbf{84.21}\stackrel{{\scriptstyle\scriptscriptstyle+32\%}}{{>}}52.63$ $\mathbf{84.91}\stackrel{{\scriptstyle\scriptscriptstyle+33\%}}{{>}}52.28$
Llama-3.2-1B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+34\%}}{{<}}\mathbf{33.68}$ $25.61\stackrel{{\scriptstyle\scriptscriptstyle+37\%}}{{<}}\mathbf{61.52}$ $\mathbf{87.25}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}65.38$ $\mathbf{90.06}\stackrel{{\scriptstyle\scriptscriptstyle+24\%}}{{>}}66.32$ $\mathbf{90.06}\stackrel{{\scriptstyle\scriptscriptstyle+25\%}}{{>}}66.08$
Llama-3.2-3B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+36\%}}{{<}}\mathbf{35.44}$ $68.19\stackrel{{\scriptstyle\scriptscriptstyle+5\%}}{{<}}\mathbf{73.10}$ $\mathbf{90.06}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}76.49$ $\mathbf{89.12}\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{>}}76.02$ $\mathbf{89.94}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}76.26$
Qwen3-4B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+54\%}}{{<}}\mathbf{47.60}$ $29.01\stackrel{{\scriptstyle\scriptscriptstyle+33\%}}{{<}}\mathbf{58.01}$ $\mathbf{84.09}\stackrel{{\scriptstyle\scriptscriptstyle+28\%}}{{>}}59.88$ $\mathbf{85.96}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}59.06$ $\mathbf{85.38}\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{>}}59.88$
Llama-3.1-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+61\%}}{{<}}\mathbf{60.58}$ $\mathbf{77.89}\stackrel{{\scriptstyle\scriptscriptstyle+2\%}}{{>}}75.79$ $\mathbf{95.67}\stackrel{{\scriptstyle\scriptscriptstyle+19\%}}{{>}}77.19$ $\mathbf{96.96}\stackrel{{\scriptstyle\scriptscriptstyle+20\%}}{{>}}77.31$ $\mathbf{97.43}\stackrel{{\scriptstyle\scriptscriptstyle+21\%}}{{>}}77.08$
Qwen3-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+62\%}}{{<}}\mathbf{58.48}$ $54.39\stackrel{{\scriptstyle\scriptscriptstyle+17\%}}{{<}}\mathbf{70.41}$ $\mathbf{95.32}\stackrel{{\scriptstyle\scriptscriptstyle+25\%}}{{>}}71.35$ $\mathbf{96.26}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}70.76$ $\mathbf{96.14}\stackrel{{\scriptstyle\scriptscriptstyle+26\%}}{{>}}71.11$
Qwen3-32B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+60\%}}{{<}}\mathbf{47.37}$ $56.14\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{<}}\mathbf{61.75}$ $\mathbf{80.12}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}62.46$ $\mathbf{80.47}\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{>}}62.34$ $\mathbf{79.77}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}62.57$
@@
Qwen3-14B $0.32\stackrel{{\scriptstyle\scriptscriptstyle+45\%}}{{<}}\mathbf{0.69}$ $\mathbf{0.79}\stackrel{{\scriptstyle\scriptscriptstyle+8\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$
@@ 5 Conclusion
Large scale language models increasingly spend the majority of their inference footprint on the KV cache, making cache compression a decisive factor in practical deployment. This work supplied a principled answer to the bit-allocation problem by unifying theory and experiment.
  • Key–Value Norm Disparity. We proved that Key weights exhibit larger Frobenius norms than Value weights throughout training, implying a higher information density in the key path.
  • Key-Driven Quantization. Leveraging this norm gap, a second theorem shows that assigning the finer bit-width to keys and the coarser bit-width to values strictly minimises quantization error, thereby preserving downstream accuracy.
  • Empirical Validation. Experiments across nine model families, three generative benchmarks, and two hardware-aligned back-ends confirm the theory: a $4\!\times\!2$ (keys / values) split consistently recovers $\geq 94\%$ of full-precision accuracy while halving KV-cache memory.
These findings elevate bit allocation from a hyper-parameter tuned by grid search to a design principle derived from model geometry, offering immediate guidance for practitioners and a foundation for hardware–algorithm co-design. We used a context length of 2K tokens, a common setting, but will use longer contexts in future.
@@ Appendix A Dynamics of the Norms of Key and Value Weight Matrices
A.1 Preliminary: Dimensions
$\begin{array}[]{@{}ll@{}}c&\text{sequence length (context size)},\\[4.0pt] d_{m}&\text{embedding dimension},\\[4.0pt] d_{v}=d_{k}=d_{m}/n_{h}=d_{m}&\text{for a single head ($n_{h}=1$)},\\[4.0pt] n_{h}&\text{number of attention heads}.\end{array}$
A.2 Symbols
$\begin{array}[]{@{}ll@{}}X&\text{input matrix }(c\times d_{m}),\\ X^{O}&\text{output matrix }(c\times d_{m}),\\ W^{V}&\text{value weights }(d_{m}\times d_{m}),\\ W^{K}&\text{key weights }(d_{m}\times d_{m}),\\ W^{Q}&\text{query weights }(d_{m}\times d_{m}),\\ K=XW^{K}&(c\times d_{m}),\\ Q=XW^{Q}&(c\times d_{m}),\\ V=XW^{V}&(c\times d_{m}),\\ S=QK^{\top}/\sqrt{d_{m}}&(c\times c),\\ A=\operatorname{softmax}(S)&(c\times c),\\ \mathcal{L}&\text{loss function.}\end{array}$
Appendix B Norm Dynamics
In this appendix, we compare the long–time behavior of the Frobenius norms of the key ($W^{K}$) and value ($W^{V}$) weight matrices of a single–head self–attention layer trained with stochastic gradient descent (SGD). We show that, under standard isotropic assumptions, $\|W^{K}\|_{F}$ grows faster than $\|W^{V}\|_{F}$.
B.1 Update rule
At update step $i$ the weights follow
$W^{m}_{i+1}\;=\;W^{m}_{i}-\eta\,\frac{\partial\mathcal{L}}{\partial W^{m}_{i}},\qquad m\in\{K,V\},$ (2)
@@ B.1 Update rule
$\|W^{m}_{i+1}\|_{F}^{2}$ $=\|W^{m}_{i}\|_{F}^{2}+\eta^{2}\Bigl{\|}\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr{\|}_{F}^{2}-2\eta\,\Bigl{\langle}W^{m}_{i},\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr{\rangle}_{F},$ (3)
@@
$\|A\|_{F}$ $\equiv\sqrt{\operatorname{tr}(A^{\top}A)},$ (4)
$\langle A,B\rangle_{F}$ $\equiv\operatorname{tr}(A^{\top}B).$ (5)
B.2 Expectation over mini-batches
Taking an expectation over mini-batches yields
$\Delta\mathbb{E}\bigl{[}\|W^{m}_{i}\|_{F}^{2}\bigr{]}$ $\equiv\mathbb{E}\bigl{[}\|W^{m}_{i+1}\|_{F}^{2}\bigr{]}-\mathbb{E}\bigl{[}\|W^{m}_{i}\|_{F}^{2}\bigr{]}$ (6)
[…]
In high–dimensional weight space the gradient is almost orthogonal to the current weights, making the second term negligible. Hence,
$\Delta\mathbb{E}\bigl{[}\|W^{m}_{i}\|_{F}^{2}\bigr{]}\;\approx\;\eta^{2}\,\mathbb{E}\Bigl{[}\Bigl{\|}\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr{\|}_{F}^{2}\Bigr{]}.$ (8)
B.3 Initialization
Throughout we assume Xavier initialization: each entry of $W^{K}$ and $W^{V}$ is drawn i.i.d. from a zero–mean distribution whose variance preserves the input scale [36].
B.4 Gradients
@@
$d\mathcal{L}(X^{O})$ $=\operatorname{tr}\!\bigl{[}(\partial\mathcal{L}/\partial X^{O})^{\!\top}\,AX\,dW^{V}W^{O}\bigr{]}=\operatorname{tr}\!\bigl{[}(AX)^{\top}\,\frac{\partial\mathcal{L}}{\partial X^{O}}\,(W^{O})^{\top}\,dW^{V}\bigr{]},$ (10)
@@
$\frac{\partial\mathcal{L}}{\partial W^{V}}=(AX)^{\top}\,\frac{\partial\mathcal{L}}{\partial X^{O}}\,(W^{O})^{\top}.$ (11)
@@
$d\mathcal{L}(S)$ $=\operatorname{tr}\!\bigl{[}(\partial\mathcal{L}/\partial S)^{\!\top}\,dS\bigr{]}=\frac{1}{\sqrt{d_{m}}}\,\operatorname{tr}\!\bigl{[}X^{\top}\,(\partial\mathcal{L}/\partial S)^{\!\top}\,Q\,dW^{K}\bigr{]},$ (12)
@@
$\frac{\partial\mathcal{L}}{\partial W^{K}}=\frac{1}{\sqrt{d_{m}}}\,X^{\top}\,(\partial\mathcal{L}/\partial S)^{\!\top}\,Q.$ (13)
B.5 Expected gradient norms
Squaring the Frobenius norm of (11) and taking an expectation under the isotropic-input assumption,
$\mathbb{E}\!\bigl{[}\|\partial\mathcal{L}/\partial W^{V}\|^{2}_{F}\bigr{]}$ $=c\,\sigma_{x}^{2}\,\sigma_{o}^{2},$ (14)
@@ B.5 Expected gradient norms
$\mathbb{E}\!\bigl{[}\|\partial\mathcal{L}/\partial W^{K}\|^{2}_{F}\bigr{]}=\frac{\sigma_{x}^{2}\,\sigma_{s}^{2}}{d_{m}}\,\|Q\|_{F}^{2},$ (15)
@@
B.6 Asymptotic comparison
Substituting these expectations into (8) shows that
$\mathbb{E}\!\bigl{[}\|W^{K}\|^{2}_{F}\bigr{]}\text{grows as}\;\eta^{2}\,\frac{\sigma_{x}^{2}\sigma_{s}^{2}}{d_{m}}\,\|Q\|_{F}^{2},\quad\text{whereas}\quad\mathbb{E}\!\bigl{[}\|W^{V}\|^{2}_{F}\bigr{]}\;\text{grows as}\;\eta^{2}\,c\sigma_{x}^{2}\sigma_{o}^{2}.$ (16)
@@ B.6 Asymptotic comparison
$\mathbb{E}\!\bigl{[}\|W^{K}\|_{F}\bigr{]}\;>\;\mathbb{E}\!\bigl{[}\|W^{V}\|_{F}\bigr{]}.$ (17)
As shown in Equation 17, the expectation of the Frobenius norm of $W^{k}$ is greater than that of $W^{v}$. [^3] While this analysis assumes a single-head, decoder-only transformer for clarity, the extension to multi-head attention is direct. The same conclusion also holds for group-query and multi-query attention.
Appendix C Quantization and Dequantization Error Bounds
In this appendix we establishes upper bounds on the quantization error incurred when a matrix is represented using a finite bit-width in two’s-complement format. Two theorems are presented: one that characterizes the error in terms of the spectral norm and another in terms of the Frobenius norm.

### v3

@@ 4.3 Mixed-Precision Downstream Accuracy
Qwen3-0.6B $00\stackrel{{\scriptstyle\scriptscriptstyle+3\%}}{{<}}\mathbf{0.03} 0.0$$0\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{<}}\mathbf{0.16} 0.0$$04\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{<}}\mathbf{0.17}$ $\mathbf{0.34}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.18$ $\mathbf{0.34}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.18$
Llama-3.2-1B $00\stackrel{{\scriptstyle\scriptscriptstyle+5\%}}{{<}}\mathbf{0.05} 0.0$$01\stackrel{{\scriptstyle\scriptscriptstyle+18\%}}{{<}}\mathbf{0.19}$ $\mathbf{0.27}\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{>}}0.20$ $\mathbf{0.28}\stackrel{{\scriptstyle\scriptscriptstyle+8\%}}{{>}}0.20$ $\mathbf{0.28}\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{>}}0.21$
@@
Qwen3-4B $00\stackrel{{\scriptstyle\scriptscriptstyle+41\%}}{{<}}\mathbf{0.41} 0.0$$01\stackrel{{\scriptstyle\scriptscriptstyle+49\%}}{{<}}\mathbf{0.50}$ $\mathbf{0.80}\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{>}}0.51$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.51$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.51$
@@
Qwen3-8B $00\stackrel{{\scriptstyle\scriptscriptstyle+48\%}}{{<}}\mathbf{0.48} 0.0$$08\stackrel{{\scriptstyle\scriptscriptstyle+46\%}}{{<}}\mathbf{0.54}$ $\mathbf{0.86}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.55$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+32\%}}{{>}}0.55$ $\mathbf{0.86}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.55$
@@
Qwen3-0.6B $00\stackrel{{\scriptstyle\scriptscriptstyle+10\%}}{{<}}\mathbf{0.10} 0.0$$00\stackrel{{\scriptstyle\scriptscriptstyle+53\%}}{{<}}\mathbf{0.53}$ $0.51\stackrel{{\scriptstyle\scriptscriptstyle+2\%}}{{<}}\mathbf{0.53}$ $\mathbf{0.84}\stackrel{{\scriptstyle\scriptscriptstyle+32\%}}{{>}}0.53$ $\mathbf{0.85}\stackrel{{\scriptstyle\scriptscriptstyle+33\%}}{{>}}0.52$
Llama-3.2-1B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+34\%}}{{<}}\mathbf{0.34}$ $0.26\stackrel{{\scriptstyle\scriptscriptstyle+37\%}}{{<}}\mathbf{0.62}$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}0.65$ $\mathbf{0.90}\stackrel{{\scriptstyle\scriptscriptstyle+24\%}}{{>}}0.66$ $\mathbf{0.90}\stackrel{{\scriptstyle\scriptscriptstyle+25\%}}{{>}}0.66$
Llama-3.2-3B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+36\%}}{{<}}\mathbf{0.35}$ $0.68\stackrel{{\scriptstyle\scriptscriptstyle+5\%}}{{<}}\mathbf{0.73}$ $\mathbf{0.90}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.76$ $\mathbf{0.89}\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{>}}0.76$ $\mathbf{0.90}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.76$
Qwen3-4B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+54\%}}{{<}}\mathbf{0.48}$ $0.29\stackrel{{\scriptstyle\scriptscriptstyle+33\%}}{{<}}\mathbf{0.58}$ $\mathbf{0.84}\stackrel{{\scriptstyle\scriptscriptstyle+28\%}}{{>}}0.60$ $\mathbf{0.86}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.59$ $\mathbf{0.85}\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{>}}0.60$
Llama-3.1-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+61\%}}{{<}}\mathbf{0.61}$ $\mathbf{0.78}\stackrel{{\scriptstyle\scriptscriptstyle+2\%}}{{>}}0.76$ $\mathbf{0.96}\stackrel{{\scriptstyle\scriptscriptstyle+19\%}}{{>}}0.77$ $\mathbf{0.97}\stackrel{{\scriptstyle\scriptscriptstyle+20\%}}{{>}}0.77$ $\mathbf{0.97}\stackrel{{\scriptstyle\scriptscriptstyle+21\%}}{{>}}0.77$
Qwen3-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+62\%}}{{<}}\mathbf{0.58}$ $0.54\stackrel{{\scriptstyle\scriptscriptstyle+17\%}}{{<}}\mathbf{0.70}$ $\mathbf{0.95}\stackrel{{\scriptstyle\scriptscriptstyle+25\%}}{{>}}0.71$ $\mathbf{0.96}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}0.71$ $\mathbf{0.96}\stackrel{{\scriptstyle\scriptscriptstyle+26\%}}{{>}}0.71$
Qwen3-32B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+60\%}}{{<}}\mathbf{0.47}$ $0.56\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{<}}\mathbf{0.62}$ $\mathbf{0.80}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}0.62$ $\mathbf{0.80}\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{>}}0.62$ $\mathbf{0.80}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}0.63$
@@
The HQQ results, shown in Table 3, extend these observations to a broader range of bit-widths, model scales, and tasks. The analysis systematically compares $\mathbf{K}_{i}\mathbf{V}_{x}$ against $\mathbf{K}_{x}\mathbf{V}_{i}$ for $i\in\{1,2,4,6,8\}$, where $x$ denotes the mean accuracy over all bit-widths of the other cache. This directly addresses the question: “If $i$ bits are available, should they be allocated to keys or values?” Across all models (0.6B-32B) and datasets (GSM8K, CoQA, EQ-Parseable), the answer is consistently “keys”. At ultra-low precision (1-2 bits), the advantage is especially pronounced; for instance, on GSM8K, allocating a single extra bit to keys with 1-bit quantization yields gains of +48 pp for Qwen3-8B. Similar trends hold on EQ-Bench and CoQA: for Qwen3-0.6B on EQ-Parseable, prioritizing keys delivers up to +62 pp, and key-first allocations never underperform value-first ones in any configuration. Even at moderate precisions (4-6 bits), key-centric allocation continues to offer 7-12 pp improvements for 8-14B models, indicating that the advantage persists well beyond extreme compression. On average, $\mathbf{K}_{4}\mathbf{V}_{2}$ retains 98.3% accuracy of $\mathbf{K}_{4}\mathbf{V}_{4}$ (CoQA: 99.2%, EQ-Bench: 99.35%, GSM8K: 97.7%; worst: 88.3%, best: 103.5%).
Figure 4: Integration of rotation and mixed-precision quantization. Downstream accuracy is shown for four quantization configurations ($\mathrm{K}_{2}\mathrm{V}_{2}$, $\mathrm{K}_{2}\mathrm{V}_{4}$, $\mathrm{K}_{4}\mathrm{V}_{2}$, $\mathrm{K}_{4}\mathrm{V}_{4}$) combined with four rotation strategies (none, key-only, value-only, both), using a fixed group size of 64 for both keys and values. Results are reported on CoQA, GSM8K, EQ-Bench, and LongBench, enabling a controlled comparison of precision-rotation interactions.
More detailed downstream accuracy results are provided in Appendix C.5. Overall, downstream accuracy is far more sensitive to key precision than to value precision. Across both token-wise and channel-wise quantization schemes, model scales, and task types, assigning the higher bit-width to $\mathrm{K}$ consistently yields near-baseline accuracy while substantially reducing KV memory. This establishes a simple, backend-agnostic design principle for mixed-precision KV cache quantization: More for keys, less for values.
4.4 Integrating with Rotation-Based Methods
Figure 4 visualizes how rotation interacts with key-value bit allocations. Across all tasks, applying rotation to keys consistently yields larger gains than applying it to values. Notably, applying key-only rotation to the $\mathrm{K}_{4}\mathrm{V}_{2}$ configuration achieves accuracy that closely matches the full $\mathrm{K}_{4}\mathrm{V}_{4}$ baseline, indicating that the primary benefits of rotation arise from mitigating key outliers. These trends are consistent across tasks and model scales, reinforcing that rotation is most effective when applied in conjunction with key-favored bit allocation.
Appendix C.6 presents detailed downstream results illustrating the synergistic effects of integrating rotation with mixed-precision quantization across tasks and models. It also examines the impact of key and value group sizes, showing that smaller sizes are beneficial for $\mathrm{K}$ due to higher information density, whereas larger group sizes suffice for $\mathrm{V}$ given lower sensitivity to quantization.
@@ 5 Conclusion
As large language models increasingly devote most of their inference cost to KV-cache storage and access, effective cache compression has become essential for practical deployment. This work provides a theoretically grounded justification and solution to the bit-allocation problem by bridging model geometry and quantization design. We theoretically establish that key projections consistently carry higher information density than value projections. Building on this, we show that allocating higher precision to keys and lower precision to values minimizes quantization error. Extensive experiments across nine model families, six benchmarks, and two hardware-aligned backends validate this principle: a $\mathrm{K}_{4}\mathrm{V}_{2}$ precision split reliably recovers up to $98.3\%$ of $\mathrm{K}_{4}\mathrm{V}_{4}$ accuracy while significantly reducing memory consumption. Moreover, we demonstrate that our geometry-driven strategy is orthogonal to rotation-based outlier redistribution methods, enabling seamless integration and further accuracy gains. These findings elevate bit allocation from empirical tuning to a theoretically grounded geometry-driven design principle, providing clear guidance for efficient deployment and future hardware-algorithm co-design.
Limitations
While our geometry-driven mixed-precision quantization framework demonstrates both theoretical soundness and practical effectiveness through rigorous analysis and extensive experiments, several limitations remain. First, all evaluations are conducted with a maximum context length of 2,000 tokens, which reflects common inference-time configurations but does not fully capture the behavior of models operating at much larger context windows. Extending the approach to longer contexts may expose additional challenges, including increased quantization sensitivity and higher memory management overheads. Addressing these factors is an important direction for future work.
Ethical Considerations
This research aims to reduce the memory and computational costs of large language model inference by utilizing efficient KV-cache quantization. Such improvements have the potential to lower energy consumption and broaden access to language models in resource-constrained settings, promoting more sustainable and inclusive deployment. However, any compression technique entails accuracy trade-offs, which must be carefully monitored to avoid disproportionate impacts in high-stakes domains such as healthcare, law, or finance. Responsible deployment requires thorough evaluation of model behavior under mixed-precision settings, particularly for safety-critical applications.
Acknowledgment
This research was supported in part by NSF awards 2117439, 2112606, and 2320952.
@@ Appendix A Dynamics of the Norms of Key and Value Weight Matrices
A.1 Preliminaries and Notation
$c$ Sequence length
$d_{m}$ Embedding dimension
$d_{v}=d_{k}=d_{m}/n_{h}$ Single head dimension
$n_{h}$ Number of attention heads
[…]
A.2 Training Dynamics of Frobenius Norms
We compare the long-time behavior of the Frobenius norms of the key ($W^{K}$) and value ($W^{V}$) weight matrices of a single-head self-attention layer trained with stochastic gradient descent (SGD). We show that, under standard isotropic assumptions, $\|W^{K}\|_{F}$ grows faster than $\|W^{V}\|_{F}$.
Update rule. At step $i$, the weights follow
$W^{m}_{i+1}\;=\;W^{m}_{i}-\eta\,\frac{\partial\mathcal{L}}{\partial W^{m}_{i}},\qquad m\in\{K,V\},$
@@ A.2 Training Dynamics of Frobenius Norms
$\|W^{m}_{i+1}\|_{F}^{2}$ $=\|W^{m}_{i}\|_{F}^{2}+\eta^{2}\Bigl\|\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr\|_{F}^{2}$
$\quad-2\eta\,\Bigl\langle W^{m}_{i},\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr\rangle_{F}.$
@@
$\|A\|_{F}$ $\equiv\sqrt{\operatorname{tr}(A^{\top}A)},$
$\langle A,B\rangle_{F}$ $\equiv\operatorname{tr}(A^{\top}B).$
Expectation over mini-batches. Taking an expectation over mini-batches yields
$\Delta\mathbb{E}\bigl[\|W^{m}_{i}\|_{F}^{2}\bigr]$ $\equiv\mathbb{E}\bigl[\|W^{m}_{i+1}\|_{F}^{2}\bigr]-\mathbb{E}\bigl[\|W^{m}_{i}\|_{F}^{2}\bigr]$
$=\eta^{2}\,\mathbb{E}\!\Bigl[\Bigl\|\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr\|_{F}^{2}\Bigr]$
$\quad-2\eta\,\mathbb{E}\!\Bigl[\bigl\langle W^{m}_{i},\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\bigr\rangle_{F}\Bigr].$
In high-dimensional weight space, the gradient is almost orthogonal to the current weights, making the second term negligible. Hence,
$\Delta\mathbb{E}\bigl[\|W^{m}_{i}\|_{F}^{2}\bigr]\;\approx\;\eta^{2}\,\mathbb{E}\Bigl[\Bigl\|\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr\|_{F}^{2}\Bigr].$
Throughout, we assume Xavier initialization: each entry of $W^{K}$ and $W^{V}$ is drawn i.i.d. from a zero-mean distribution whose variance preserves the input scale Glorot and Bengio (2010).
@@
$d\mathcal{L}(X^{O})$ $=\operatorname{tr}\!\bigl[(\partial\mathcal{L}/\partial X^{O})^{\!\top}\,AX\,dW^{V}W^{O}\bigr]$
$=\operatorname{tr}\!\bigl[(AX)^{\top}\,\frac{\partial\mathcal{L}}{\partial X^{O}}\,(W^{O})^{\top}\,dW^{V}\bigr].$
@@
$\frac{\partial\mathcal{L}}{\partial W^{V}}=(AX)^{\top}\,\frac{\partial\mathcal{L}}{\partial X^{O}}\,(W^{O})^{\top}.$
@@
$d\mathcal{L}(S)$ $=\operatorname{tr}\!\bigl[(\partial\mathcal{L}/\partial S)^{\!\top}\,dS\bigr]$
$=\frac{1}{\sqrt{d_{m}}}\,\operatorname{tr}\!\bigl[X^{\top}\,(\partial\mathcal{L}/\partial S)^{\!\top}\,Q\,dW^{K}\bigr].$
@@
$\frac{\partial\mathcal{L}}{\partial W^{K}}=\frac{1}{\sqrt{d_{m}}}\,X^{\top}\,(\partial\mathcal{L}/\partial S)^{\!\top}\,Q.$
Squaring the Frobenius norm of (A.2) and taking an expectation under the isotropic-input assumption,
$\mathbb{E}\!\bigl[\|\partial\mathcal{L}/\partial W^{V}\|^{2}_{F}\bigr]$ $=c\,\sigma_{x}^{2}\,\sigma_{o}^{2},$
@@
$\mathbb{E}\!\bigl[\|\partial\mathcal{L}/\partial W^{K}\|^{2}_{F}\bigr]=\frac{\sigma_{x}^{2}\,\sigma_{s}^{2}}{d_{m}}\,\|Q\|_{F}^{2},$
@@
Substituting these expectations into (A.2) shows that
$\mathbb{E}\!\bigl[\|W^{K}\|^{2}_{F}\bigr]$ grows as $\eta^{2}\,\frac{\sigma_{x}^{2}\sigma_{s}^{2}}{d_{m}}\,\|Q\|_{F}^{2}$, whereas $\mathbb{E}\!\bigl[\|W^{V}\|^{2}_{F}\bigr]$ grows as $\eta^{2}\,c\sigma_{x}^{2}\sigma_{o}^{2}.$
@@
$\mathbb{E}\!\bigl[\|W^{K}\|_{F}\bigr]\;>\;\mathbb{E}\!\bigl[\|W^{V}\|_{F}\bigr].$
As shown in Equation A.2, the expectation of the Frobenius norm of $W^{k}$ is greater than that of $W^{v}$. [^2] The analysis considers a single-head decoder-only Transformer without loss of generality; the conclusions extend directly to multi-head, group-query, and multi-query attention.
Appendix B Quantization Error Bounds
We establish upper bounds on the quantization error incurred when a matrix is represented using a finite bit-width in two’s complement format. Two theorems are presented: one that characterizes the error in terms of the spectral norm and another in terms of the Frobenius norm.